Shift report 20030202 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Feb 02 2003 - 16:39:20 EST

  • Next message: Ian Bearden: "Re: Shift report 20030202 08:00-16:00"
    ------------------------------------------------------------
    Date: Sun Feb  2 16:39:20 2003, Shift: 08:00-16:00
    Supervisor: kh, with: zy
    Log book#13 pages 140 - 142
    
    Shift Summary:
    ------------------------------------------------------------
    Beam during the entire shift (until 13:45)
    
    Runs taken:
    Run     FS Setting  trig 6  MRS Setting  trig 3
    7737      1/8A,      181K       6A        188K
    7738      1/4A,       35K       6A         19K
    7739      1/4A,        1K       6A          3K  T1 tripped
    7740      1/4A,       59K       6A        106K
    7741      1/4A,        6K       6A         11K
    7742      1/4A,       24K       2A         58K
    
    Beam dumped at 13:35
    
    As noted above, T1 tripped.  RD asked for an access to 
    unground T1.  He couldn't do it between this store.  MCR
    said maybe after next store.  If this coming store is as
    good as the last, that would be late at night.
    
    T5 tripped at the end; probably when they dumped the beam.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun Feb 02 2003 - 16:43:25 EST